Determine the truth value of each of these statements if the domain of each variable consists of all real numbers.


a. 3x (x ² = 2)


b. 3x (x ² = -1)


c. Vx (x ²+2 à ¢ ‰ ¥ 1)


d. Vx (x ² + x)
1
Expert's answer
2021-11-02T19:02:17-0400

Part a

"\exist x (x^2 =2)\\
Here \space x= ∓ \sqrt{2} \space \space so \space satisfy". Ture


Part b

"\exist x (x^2 =1)\\"

Since for no real number, the square can be negative. So it is false.


Part c

"\forall x (x^2+2*)"

* These symbols are not proper of the reformulate up to the best of my knowledge "\forall x (x^2 \land x \not= 1)" False


Part d

"\forall x (x^2+x)"

Since for every real number x2+x has a solution. True